OpenStudy (anonymous):

(4x-3) --------------------- (24x^(3)-30x^(2)-29x-15) ---means to divide

OpenStudy (anonymous):

Actually this is a long division question

OpenStudy (anonymous):

so the problem is to divide (4x-3) by (24x^(3)-30x^(2)-29x-15) ? Is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you sure its not the other way around?

OpenStudy (anonymous):

well in any case. note that \[x = \frac{1}{\frac{1}{x}}\]

OpenStudy (anonymous):

so, your expression can be written as \[1 \div \frac{24x^3 - 30 x^2 - 29x -15}{4x-3}\]

OpenStudy (anonymous):

now just use long division to get the answer.